### v2

# A Foundational Individual Mobility Prediction Model based on Open-Source Large Language Models

Abstract
Individual mobility prediction plays a key role in urban transport, enabling personalized service recommendations and effective travel management. It is widely modeled by data-driven methods such as machine learning, deep learning, as well as classical econometric methods to capture key features of mobility patterns. However, such methods are hindered in promoting further transferability and generalization due to limited capacity to learn mobility patterns from different data sources, predict in out-of-distribution settings (a.k.a “zero-shot”). To address the challenge, this paper proposes the foundational MoBLLM model based on lightweight open-source large language models (LLMs) fine-tuned on mobility datasets. We conduct extensive experiments on six real-world mobility datasets to evaluate its performance and transferability for diverse prediction tasks across varying times (years), space (cities), and situations (e.g., disruptions). The MoBLLM model achieves best F1 and accuracy performance in all datasets Compared with the state-of-the-art deep learning models. It also shows better transferability and cost performance than commercial LLMs. Further experiment reveals its high adaptability to different situations under network change, policy intervention, special events, and incidents. These results demonstrate the foundational capability of the MoBLLM model in predicting individual mobility, which contributes to precise and robust provision of personalized information for transportation management.
keywords:
Individual Mobility Prediction , Large Language Models , Fine-Tuning , Foundational Model
journal: Transportation Research Part C
[…]
[label1]organization=Department of Civil and Architectural Engineering, KTH Royal Institute of Technology,country=Sweden
[…]
[label2]organization=Department of Data Science and Artificial Intelligence, Monash University,country=Australia
[…]
[label3]organization=Department of Technology, Management and Economics Intelligent Transportation Systems, Technical University of Denmark,country=Denmark
1 Introduction
Individual mobility prediction studies human mobility patterns at an individual level to forecast future movements. It facilitates a broad spectrum of applications in transportation domain, such as personalized services of POI or route recommendations and incentive mobility management (Ma and Zhang, 2022). Data-driven methods are dominant in this field including conventional Markov-based machine learning models (Gambs et al., 2012; Mathew et al., 2012; Lu et al., 2013) and deep learning (DL) models such as recurrent neural network (RNN) or Transformer (Feng et al., 2018; Kong and Wu, 2018; Xue et al., 2021). However, they are still limited to specialized training on the mobility dataset from a single city or data source, leading to low transferability of adapting to different travel contexts in different cities. It would be valuable to develop a foundational individual mobility prediction model capable of addressing these challenges.
Recently, Large Language Models (LLMs) have made significant progress, far surpassing previous state-of-the-art models such as BERT (Devlin et al., 2019) and T5 (Raffel et al., 2020). The LLMs become the new foundational model in Natural Language Processing (NLP) domain. In Artificial Intelligence (AI) systems, a foundational model is any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks (Bommasani et al., 2022). As fine-tuning techniques advance, LLMs can be applied to developing foundational models for other various disciplines (Wang et al., 2025; Awais et al., 2025). Inspired by the concept of foundational models in AI, we define the foundational mobility prediction model as one trained on extensive mobility data, capable of adapting to diverse prediction tasks across varying time scales (years), spatial scales (cities), and situational conditions (such as disruptions and interventions).
LLMs have also inspired a range of applications in urban mobility (Ma et al., 2025; Wang et al., 2024), particularly in mobility prediction tasks. Specifically, typical LLM based methods like In-Context Learning (ICL) (Brown et al., 2020) and Chain-of-Thought (CoT) (Wang et al., 2023a; Wei et al., 2022) enable modeling mobility prediction problems with natural language (Wang et al., 2023b; Liang et al., 2024; Feng et al., 2024a). However, hallucination issues (Huang et al., 2025) may arise when applying these methods, as their use of commercial LLMs lacks specialized training on mobility datasets. Therefore, the existing prompt engineering based methods are limited in generalizing prediction under different contexts (lack foundational capability). Besides, fine-tuning based on pretrained LLMs are applied to developing the individual mobility prediction framework (Li et al., 2024; Gong et al., 2024). They begin by aligning visited locations with specific intents, then leverage LLMs to predict the next location based on this enriched semantic context. From model validation perspectives, most of studies focus on prediction / accuracy but have not sufficiently demonstrated the LLMs’ capabilities of generalization and transferability across different data sources and prediction scenarios.
To fill the gap, we propose a foundational individual mobility prediction model MoBLLM based on open-source LLMs. It uses the pre-trained LLMs as the core module with generic prediction prompts and introduces instruction tuning (Zhang et al., 2023c) to enhance the LLMs’ capacities in adapting to different prompt writing styles of users and ensuring high transferability on different types of mobility data (e.g., check-in data). It adopts the Parameter Efficient Fine-Tuning (PEFT) method to realize the parameter and cost efficiency of training compared with the conventional full-parameter fine-tuning method. The key contributions include:
  • 1.
    Propose a foundational, open-source LLM-based individual mobility prediction model, which can adapt to different data from multiple sources and cities, various individual mobility prediction tasks and even diverse prediction scenarios.
  • 2.
    Introduce an efficient instruction generation approach to produce plentiful CoT based prompts for adapting to different writing styles of users.
  • 3.
    Validate the proposed method through extensive experiments on different mobility datasets, including GPS location, AFC, and check-in data from different cities over the world. It compares against state-of-the-art DL and LLM models in aspects of prediction accuracy and transferability (space, time, and situation).
The remainder of the paper is organized as follows: Section 2 reviews the recent approaches related to modeling individual mobility prediction in the literature. Section 3 defines the studied problem and proposes the framework, including instruction data generation and fine-tuning based on open-source LLMs. Case studies are conducted on real-world datasets are presented in Section 4. The final section concludes the main findings and discusses future work.
2 Related Work
The problem of individual mobility prediction consists of predicting an individual’s possible next location or trip information given their historical mobility sequence data (Ma and Zhang, 2022). In early stage, the Markov based methods are widely applied to modeling location transition probabilities, i.e, the next location prediction is based on several previous visited locations (Gambs et al., 2012; Mathew et al., 2012; Lu et al., 2013). Although these methods are intuitive and interpretable, they have low performance due to the difficulty in learning complex mobility patterns. Therefore, we focus on literature based on deep learning and LLM.
2.1 Deep Learning based Methods
Deep learning based methods commonly model individual mobility prediction by capturing the spatiotemporal features of an individual’s historical trip sequence. For example, Liu et al. (2016) models local temporal and spatial contexts by extending an recurrent neural network (RNN) layer with time-specific and distance-specific transition matrices. DeepMove (Feng et al., 2018) adopted a sequence to sequence (seq2seq)-based framework to capture the multi-level periodicity for mobility prediction from lengthy and sparse trajectories. The DeepTrip model (Zhang et al., 2023a) extended the DeepMove to predict the next trip information of the AFC metro data. Other RNN-based models adopted the Long Short-Term Memory (LSTM) architecture to model the prediction problem (Kong and Wu, 2018; Li et al., 2020; Chen et al., 2020).
Besides, several Transformer-based models show stronger capacities in capturing spatiotemporal context features of human trajectories and faster inference speed than the RNN-based models. Specifically, Xue et al. (2021) used a Transformer-based framework to learn temporal, semantic, social, and geographical contexts from the history of place sequences to predict the next place. Hong et al. (2023) adopted a multi-head self-attention mechanism model to infer the context of historical location visits, visit time, activity duration, and POI to infer individuals’ next location.
The above methods based on RNN or Transformer have achieved significantly better performance than the classical machine learning models like Markov but are still difficult to address the issue of low transferability. These methods can only be generalized to the samples aligned with the distribution of their training data since their relatively small parameter sizes limit the capability of learning the representations of mobility patterns from different data sources.
2.2 Large Language Model based Methods
Common LLM-based methods for mobility prediction employ In-Context Learning (ICL) approaches to prompt LLMs with prediction task descriptions and historical mobility data, as well as Chain-of-Thought (CoT) to guide LLMs to capture complex semantic information. Such approaches enable LLM to make predictions in a zero-shot way without requiring additional training or gradient update. For example, LLM-Mob (Wang et al., 2023b) adopted ICL to realize zero-shot individual mobility prediction. Similarly, LLM-MPE (Liang et al., 2024) predicted human mobility under public events by leveraging the unprecedented ability of LLM to process textual data. AgentMove (Feng et al., 2024a) achieved a generalized prediction of mobility for any city worldwide by introducing POI data from cities to capture the knowledge of the world. LingoTrip Qin et al. (2025) prompted LLMs to capture trip purposes, choice preferences and choice variability from individual’s historical trip data to comprehensively infer the next trip location. While these methods exhibit promising application potential, the performance still has significant room for improvement since they use commercial LLMs trained on millions of general corpora without training on any mobility dataset. Meanwhile, the commercial LLMs lacking specialized knowledge may cause hallucination issues when answering domain-specific questions (Huang et al., 2025).
To solve these problems, fine-tuning methods based on open-source pre-trained LLMs such as Llama (Touvron et al., 2023) adapt the LLMs to individual mobility datasets. Specifically, LIMP (Li et al., 2024) proposed a mixed framework that uses GPT-4o to generate high-quality intent data for training a local open-source LLM. Mobility-LLM (Gong et al., 2024) utilized LLMs to infer travel preferences using historical mobility data and POI information. Although these methods enable LLMs to capture the mobility intents or preferences of individual by fine-tuning, they have not demonstrated the capability of adapting to different data sources across multiple cities and diverse prediction scenarios such as disruptions and interventions.
Overall, the existing methods based on prompt engineering or fine-tuning show high potential in modeling the individual mobility prediction problem and yield a wide range of valuable applications. However, they are still insufficient to build a foundational model for individual mobility prediction due to the unsolved issues in adapting to diverse prediction tasks across varying spatiotemporal scales and conditional scenarios.
3 Methodology
3.1 Problem Definition
Most mobility prediction studies focus on predicting the next location given a series of historical mobility records of the individual (Ma and Zhang, 2022). Considering GPS trajectory data, check-in data, and AFC data used in these studies, we define four base tasks of individual mobility prediction tailored to different types of mobility data.
Task 1 (GPS trajectory location prediction): Given a user’s sequence of $n$ chronological stays: $\textbf{S}=(S_{1},S_{2},\dots,S_{n})$, the task is to predict the next location of $S_{n+1}$. A stay $S_{i}$ includes the start time of stay $t_{i}$, day of the week $w_{i}$, duration $dur_{i}$ of $S_{i}$, and location $l_{i}$. Note that the location $l_{i}$ denotes the corresponding GPS point in the original trajectory after discretization and the duration is $dur_{i}=t_{i}-t_{i-1}$.
Task 2 (check-in location prediction): Given a user’s sequence of $n$ chronological stays: $\textbf{S}=(S_{1},S_{2},\dots,S_{n})$, the task is to predict the next location of $S_{n+1}$. A stay $S_{i}$ includes the start time of stay $t_{i}$, day of the week $w_{i}$, and location $l_{i}$ (physical location / point of interest).
Task 3 (next trip origin prediction): Given a user’s sequence of $n$ chronological trip activities: $\textbf{S}=(S_{1},S_{2},\dots,S_{n})$, the task is to predict the next trip end location of $S_{n+1}$. A trip activity $S_{i}$ includes the start time $t_{s_{i}}$, day of the week $w_{i}$, duration $dur_{i}$ of $S_{i}$, start location $l_{s_{i}}$, and end location $l_{e_{i}}$. We define end time $t_{e_{i}}$, and then the trip activity duration is $dur_{i}=t_{e_{i}}-l_{s_{i}}$.
Task 4 (next trip destination prediction): Given a user’s sequence of $n-1$ chronological trips: $\textbf{Tr}=(Tr_{1},Tr_{2},\dots,Tr_{n-1})$, the task is to predict the next destination station of $Tr_{n}$. A trip $Tr_{i}$ includes start time of trip $t_{o_{i}}$, day of the week $w_{i}$, origin station $s_{o_{i}}$, duration $dur_{i}$ of last stay occurring before $Tr_{i}$, and destination station $d_{i}$. We define end time of trip $t_{d_{i}}$, and then the stay duration is $dur_{i}=t_{o_{i}}-t_{d_{i-1}}$.
The definitions of Tasks 1 and 2 are commonly used in studies conducted on GPS trajectory data or check-in data (Wang et al., 2023b; Li et al., 2024; Feng et al., 2024a) as shown in Figure 1. The GPS trajectory data are transformed into discrete representations due to the sparsity of their original coordinate format, making it difficult to model the semantic features of a location. For certain check-in data using the definition of Task 2, the duration information between two adjacent stays can be omitted due to a probable long time span, providing little effective activity information.
Tasks 3 and 4 are generally defined in the studies conducted on AFC trip data. Here, we adopt an activity-based prediction framework assuming that a stay between two trips contains only one activity, as shown in Figures 1 and 1, which may benefit the model performance (Mo et al., 2021; Qin et al., 2025). Specifically, we define $d_{0}=home$ and $t_{0}=4:00AM$, which means the content of the first activity can be confirmed as staying home. The activity after the last trip of the day is not considered because it is exactly staying home the next day. The information of a trip mainly includes origin station $o$, destination station $d$, start time of trip $t_{o}$, end time of trip $t_{d}$, week of day $w$, which are commonly employed in trip prediction studies (Zhang et al., 2023b; Zhao et al., 2018). However, we extend a trip with the duration of the last stay defined in Task 4, supposing it is associated with certain activity information. Given the context of prediction on AFC trip data, we can align some notations of Task 3 with the trip definition where $t_{s_{i}}=t_{d_{i}}$, $t_{e_{i}}=t_{o_{i+1}}$, $l_{s_{i}}=d_{i}$ and $l_{e_{i}}=o_{i+1}$.
Figure 1: The examples of base tasks. The subfigure (a) presents an example of Task 1 or Task 2, and Tasks 3 and 4 are respectively illustrated by subfigures (b) and (c).
3.2 MoBLLM Framework
Figure 2 shows the framework of foundational MoBLLM model, focusing on the fine-tuning pipeline that consists of three main steps. It is basically using a top-tier versatile commercial LLM as the teacher model to instruct a customized and dedicated student model based on open-source LLMs. Specifically, first it uses advanced commercial LLMs (e.g., OpenAI’s GPT series) for fine-tuning by generating multi-style semi-complete instructions based on predefined base task prompt templates. These instructions are then combined with real-world mobility data to construct complete instruction samples. Finally, PEFT methods are applied to train an open-source LLM, such as Meta’s LLaMA series (Touvron et al., 2023), on the instruction dataset for mobility prediction tasks. The resulting fine-tuned model is referred to as MoBLLM. It can handle multiple individual mobility prediction tasks by prompting necessary data context and guidance as shown in Figure 3.
Figure 2: The overview of the proposed MoBLLM framework.
3.3 Prompt Template for Individual Mobility Prediction
To prompt LLMs to predict individual mobility in a batch processing manner, we employ an efficient template based on zero-shot CoT, following similar structures commonly adopted in recent literature (Wang et al., 2023b; Qin et al., 2025; Liang et al., 2024). The zero-shot CoT enhances LLMs to capture the spatiotemporal mobility patterns from historical data. As shown in Figure 3, the prompt template includes five components: task definition, data description, thinking guidance, output format, and data inputs.
First, the template briefly introduces the prediction tasks according to the definitions of base tasks in Section 3.1. To improve LLMs’ capability of understanding the context of mobility data, we provide detailed descriptions about the formatted data, which are adjusted due to the practical data usage of the corresponding tasks. The output format regulates the LLM’s output by asking it to provide its prediction and reason, facilitating quick processing of the text output. Such an approach of asking explanation not only improves LLMs’ prediction performance in mobility prediction but also increases the interpretability and reliability of the model (Wang et al., 2023b).
The thinking guidance is the key to carrying out the LLMs’ zero-shot learning in individual mobility prediction. It guides LLMs to capture mobility patterns from different time scales, i.e., the first and second aspects of thinking guidance shown in Figure 3. Then the third aspect reminds LLMs to learn the semantic information from the trips’ spatiotemporal context.
3.4 Instruction Data Generation
Since most current LLMs are initially designed for dealing with various natural language tasks, there is a non-trivial mismatch between the original training objective of LLMs and our studied individual mobility prediction problem. To address the mismatch, we adopt the instruction tuning approach, which is commonly applied to enhance the capabilities and controllability of LLMs in other domain-specific tasks (Zhang et al., 2023c). The key to the approach is to construct an instruction-tuning dataset for LLMs’ fine-tuning. Each sample in the dataset can be concisely represented as an $(Instruction,Output)$ pair. $Instruction$ is a kind of prompt for instructing LLMs to specialize in the studied problem during the fine-tuning process, as shown in Figure 3. The $Output$ denotes the desired output that follows the $Instruction$.
In practical application scenarios, users have different styles in writing prompts. The LLMs fine-tuned by single-style instructions have difficulty adapting to varying prompts, which probably leads to model performance reduction. To address this issue, we adopt an Alpaca-style prompt (Taori et al., 2023) to distill different expressions describing individual mobility prediction tasks to generate sufficient multi-style task instructions as the examples shown in Figure 2. Typically, distillation imparts knowledge and cognitive abilities from a competent teacher model to a simpler and more computationally efficient student model (Zhang et al., 2023c).
Figure 4 shows the CoT-based prompt for generating multi-style instructions with three main parts (see Appendix for generated examples). First, the LLMs should know how to generate diverse instructions related to time series prediction. Then, the requirements assist LLMs to work like professional mobility researchers to focus on the goal of human mobility prediction. In particular, it is necessary to limit the generation scope to text and obligatorily consider the generated instruction structure in the last three requirements. Such structural information can guide LLMs to generate CoT-based task instructions, which have been demonstrated to improve LLMs’ mobility prediction performance by previous studies (Wang et al., 2023b; Liang et al., 2024; Feng et al., 2024b; Qin et al., 2025).
To ensure the data authenticity, only the task definition, data description, and thinking guidance in Figure 3 are generated with diverse styles to get semi-complete instructions in Figure 2. Then it combines with a piece of real-world mobility data from the database to get an $Instruction$. To align with the required prediction results of the task prompt, the corresponding $Output$ given an $Instruction$ is set as $\{prediction:place\_id\}$ where the $place\_id$ is the ground truth of the next location or origin / destination station defined in the Problem Definition.
Figure 3: The example of instruction data. For the data input, $<$history$>$ and $<$context$>$ are the user’s mobility sequences respectively, containing long-term and recent stays or trips. $<$target$>$ includes the information about current states except for location, e.g., the tap-out station and time of the last trip.
Figure 4: The prompt example to generate multi-style instructions. A prompt template of the base task is put at the end, prompting LLMs to learn the style.
3.5 Parameter Efficient Fine-Tuning
Fine-tuning is a widely adopted approach for adapting LLMs to domain-specific datasets. It leverages the pre-trained knowledge embedded in LLMs to enhance task-specific performance while requiring relatively limited data and computational resources (Parthasarathy et al., 2024). However, conventional approaches such as full-parameter fine-tuning demand substantial Graphics Processing Unit (GPU) memory and computational overhead, making them impractical for use on resource-constrained devices or small-scale workstations. In contrast, parameter-efficient fine-tuning (PEFT) methods significantly reduce memory requirements while maintaining high adaptation performance (Han et al., 2024). One of the most prominent PEFT techniques is Low-Rank Adaptation (LoRA) (Hu et al., 2021), which has been extensively used to fine-tune LLMs in mobility-related applications (Li et al., 2024; Gong et al., 2024).
Figure 5 shows that LoRA introduces two trainable matrices, $W_{1}\in\mathbb{R}^{d\times r}$ and $W_{2}\in\mathbb{R}^{r\times k}$, where the rank $r$ is significantly smaller than $\min(d,k)$. Given an input $X\in\mathbb{R}^{d\times k}$, the original pre-trained weight matrix $W_{0}$ remains fixed throughout the fine-tuning process. This matrix encodes general-purpose knowledge acquired during large-scale pretraining. In contrast, $W_{1}$ and $W_{2}$ are lightweight adapters that capture task-specific patterns—in this case, individual mobility behavior. The adapted representation used for downstream mobility prediction tasks is computed as:
$h=W_{0}X+\frac{\alpha}{r}W_{1}W_{2}X,$ (1)
where $h$ denotes the output embedding aligned with mobility-related semantics in $X$. Here, $\alpha$ is a tunable scaling factor used to control the contribution of the low-rank update.
Instead of updating the full weight matrix, LoRA introduces a low-rank correction term $W_{1}W_{2}$ on top of $W_{0}$, allowing the model to efficiently adapt without retraining all parameters. Since the number of parameters in $W_{1}$ and $W_{2}$ is orders of magnitude smaller than in $W_{0}$, this approach significantly reduces GPU memory requirements. Empirical studies report that LoRA can achieve up to a 3 times reduction in memory usage while maintaining competitive performance (Hu et al., 2021), making it well-suited for fine-tuning LLMs on mobility datasets under limited resources.
In the fine-tuning phase, the individual mobility prediction task is framed as a next-token generation task. With the instruction dataset constructed by Section 3.4, the PEFT method can tune an open-source LLM’s parameters by optimizing the following objective function in a supervised way:
$L(\theta)=-\mathbb{E}_{(x,y)\in\mathbb{D}}[\sum_{i=1}^{n}log_{\theta}P(y_{i}|x;y_{1:i-1})],$ (2)
where $\mathbb{D}$ denotes the instruction dataset and $\theta$ represents the trainable parameters in PEFT method (i.e., $W_{1}$ and $W_{2}$ in Equation 1). Each $(x,y)$ represents a pair of $(Instruction,Output)$ in the instruction dataset. The $y_{1:i-1}$ and $y_{i}$ respectively denote the previously generated tokens and the ground truth of the next token.
Figure 5: The framework of LoRA.
4 Case Studies
4.1 Data Description and Performance metrics
We aim to systematically validate the MoBLLM model capability in generalizing to diverse prediction tasks across varying time scales, spatial scales and conditional scenarios. To do that, we conduct case studies using six mobility datasets, i.e., Geolife  [^1] https://www.microsoft.com/en-us/research/publication/geolife-gps-trajectory-dataset-user-guide/, FSQ-NYC, FSQ-TKY (Yang et al., 2014), FSQ-Global [^2] https://sites.google.com/site/yangdingqi/home/foursquare-dataset (Yang et al., 2016), Hong Kong (HK) AFC data, Hangzhou (HZ) AFC data  [^3] https://tianchi.aliyun.com/dataset/21904. The detailed statistics of these datasets are presented in Table 1. According to Tasks 3 and 4 for predicting origin and destination, each AFC data can be processed to obtain two subsets, such as HK-ORI and HK-DEST.
Table 1: The descriptions of datasets.
Dataset Type User Number Test Samples District Task Type
Geolife GPS trajectory 45 3459 Beijing task 1
FSQ-NYC check-in 535 12141 New York task 2
FSQ-TYK check-in 450 11250 Tokyo task 2
FSQ-Global check-in 1500 43500
415 cities in 77
countries.
task 2
HK-ORI AFC 500 5260 Hong Kong task 3
HK-DEST AFC 500 5260 Hong Kong task 4
HZ-ORI AFC 1200 15296 Hangzhou task 3
HZ-DEST AFC 1200 15296 Hangzhou task 4
In the case studies, we employ the following two metrics to evaluate the predictive performance of models:
  • 1.
    Accuracy: measures if the next location’s prediction is correct compared with the ground truth of the next visited location. We use $ACC(\%)$ in the following parts to denote this metric that measures the ratio of correct predictions in the test set, which is formulated as follows:
    $ACC=\frac{T}{T+N}\times 100\%,$ (3)
    where $T$ and $N$ are the number of correctly and incorrectly predicted samples, respectively.
  • 2.
    Weighted F1 measures the model’s performance in predicting frequently visited locations to avoid the issue of imbalance of individual’s visited locations, which is denoted as the abbreviation $F1$ in the following parts.
    $F1=\frac{2TP}{2TP+FP+FN},$ (4)
    where $TP$, $FP$ and $FN$ are respectively the number of true positive, false positive and false negative samples.
4.2 PEFT methods and parameters
To adapt open-source LLMs to real-world human mobility prediction tasks under resource constraints, we apply LoRA and several advanced variants in fine-tuning and compare their efficiency. These include OLoRA (Büyükakyüz, 2024), EVA (Paischer et al., 2024), PiSSA (Meng et al., 2024), LoftQ (Li et al., 2023), LoRA+ (Hayou et al., 2024), rsLoRA (Kalajdzievski, 2023), and QLoRA (Dettmers et al., 2023).
There are two important hyperparameters in Equation (1) that have a significant impact on fine-tuning. They are the rank $r$ and scaling factor $\alpha$, which control the updated parameter size and the ratio of specific knowledge influencing the fine-tuning process. Their specific values used in the PEFT methods in Table 2 are referred to in the corresponding literature, considering the trade-off between GPU memory, fine-tuning time, and model performance. All PEFT methods used in this research are based on a pre-trained LLaMA-3.1-8B-Instruct  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct except the LoftQ starting with a 4-bit and 64-rank checkpoint of LLaMA-3-8B-Instruct  [^5] https://huggingface.co/LoftQ/Meta-Llama-3-8B-Instruct-4bit-64rank provided. About 100K training samples are used for fine-tuning and the batch size is set to 3 for all the methods. All the model training and experimental evaluations are conducted on a workstation with Intel(R) Xeon(R) w7-3465X CPU, 4 NVIDIA RTX 6000 Ada GPUs, and 512-Gb memory.
As shown in Table 2, the OLoRA achieves the best performance on average with the highest GPU memory consumption but the lowest fine-tuning time. The QLoRA has the lowest performance, though it needs the least GPU memory for fine-tuning because it represents weights and activations with lower-precision data types by a 4-bit quantization technique. For the following experiments, we adopt the checkpoint trained by OLoRA to compare the MoBLLM with other baseline models.
Table 2: The performance of MoBLLM model with different PEFT methods. The MEM denotes the GPU memory (Gb) and Time denotes the fine-tuning time (hour).
PEFT Geolife FSQ-NYC HK-ORI HK-DEST MEM Time
ACC F1 ACC F1 ACC F1 ACC F1
LoRA 58.31 0.5385 41.58 0.3687 84.52 0.8447 70.3 0.6972 50 30
OLoRA 58.51 0.5394 41.21 0.3651 87.86 0.8784 70.93 0.7044 60.4 23
EVA 57.13 0.5251 40.51 0.3568 87.43 0.8743 70.23 0.6979 54.3 25.5
PiSSA 53.89 0.4829 35.21 0.3053 62.39 0.6259 41.77 0.3972 60 32
LoftQ 58.5 0.537 40.92 0.3611 87.46 0.8745 70.48 0.6995 92.3 37.5
LoRA+ 58.14 0.5346 39.86 0.3487 63.75 0.6348 41.52 0.3861 54.7 32
rsLoRA 59.01 0.5414 41.55 0.3678 87.48 0.8749 68.56 0.6805 58.5 32
QLoRA 38.48 0.3348 18.15 0.1355 40.08 0.3883 42.7 0.4028 46.4 37.4
4.3 Baselines
To benchmark the performance of MoBLLM, we compare it with 4 baseline models as follows:
  • 1.
    DeepMove (Feng et al., 2018) uses recurrent networks with an attention mechanism to capture multi-scale temporal dependencies of mobility patterns.
  • 2.
    MobTCast (Xue et al., 2021) is based on a transformer encoder network to leverage temporal, semantic, social, and geographical contexts from historical mobility sequences to predict the next location.
  • 3.
    MHSA (Hong et al., 2023) uses a multi-head self-attentional network to infer an individual’s next location by considering historical location visits, visit time, activity duration, as well as their surrounding land use information.
  • 4.
    LLM-Mob (Wang et al., 2023b) is an ICL-based model to first use LLMs to predict future human mobility by zero-shot CoT.
All the deep learning models DeepMove, MobTCast and MHSA are respectively trained and tested on each dataset with the same sample size as the MoBLLM. The MoBLLM model is fine-tuned on a mixed dataset combined with Geolife, FSQ-NYC, HK-ORI, and HK-DEST where each sample is assembled with an individual historical mobility sequence and instruction context generated by GPT-4o mini of Azure OpenAI with temperature as 0.3. It consumes about 1.2M tokens of generated instructions at a friendly cost of 0.7 dollars. The LLM-Mob model also adopts GPT-4o mini with temperature as zero to test its performance on the datasets.
4.4 Prediction Performance Comparison
Table 3 shows the results of comparing the MoBLLM moel with baselines. The MoBLLM achieves the state-of-the-art performance over all the baseline models on all datasets. The deep learning models DeepMove, MobTcast, and MHSA can perform better only on HK-ORI and HK-DEST than the LLM-Mob model. Compared with Geolife and FSQ-NYC, which contain various activity patterns with high sparsity, these AFC datasets have dense mobility patterns like periodical commuting. These deep learning models can easily learn dense patterns in training but have difficulty capturing sparse patterns, while the MoBLLM can adapt to different types of datasets regardless of sparse or dense patterns. The results of comparing the two LLM-based methods demonstrate that open-source LLMs can obtain substantial performance improvement by PEFT fine-tuning (better than commercial ones).
Table 3: The performance comparison between the MoBLLM and baseline models
Model Geolife FSQ-NYC HK-ORI HK-DEST
ACC F1 ACC F1 ACC F1 ACC F1
DeepMove 26.10 0.1890 19.30 0.1550 82.81 0.827 65.45 0.6450
MobTCast 29.50 0.1730 20.20 0.1660 83.28 0.8324 65.49 0.6478
MHSA 31.40 0.2180 20.20 0.1490 83.55 0.8349 66.00 0.6554
LLM-Mob 45.10 0.4040 27.40 0.2210 62.81 0.6494 54.23 0.5334
MoBLLM 58.51 0.5394 41.21 0.3651 87.86 0.8784 70.93 0.7044
4.5 Transferability Performance
Transferability measures the model’s performance in adapting to data sources that are significantly different from the training data. Based on the checkpoint trained by OLoRA in Section 4.2, the MoBLLM is tested its transferability on the datasets FSQ-TYK, FSQ-Global, HZ-ORI and HZ-DEST as described in Table 1, which are unseen in the fine-tuning. Due to the inconsistency of predictable locations between training data and the transferability test data, the conventional deep learning models DeepMove, MobTCast, and MHSA are unable to be applied to other cities’ mobility data. Therefore, we only use the LLM-Mob deployed with GPT-4o mini as the baseline in this experiment. To avoid the disturbance caused by different prompts, the MoBLLM adopts the same prompt as LLM-Mob, such as the example in Figure 3.
The results are presented in Table 4, where the open-source MoBLLM model significantly outperforms the general-purpose commercial LLM. It demonstrates that an open-source LLM, i.e., LLaMA-3.1-8B-Instruct used in the case studies, can become an expert in different mobility prediction tasks by the proposed fine-tuning framework. Moreover, the MoBLLM shows a significant superiority in cost efficiency, mainly produced by generating instructions for training, which is about 60 times cheaper than a commercial LLM.
Table 4: The transferability and cost comparison between GPT-4o mini and MoBLLM on four datasets. The cost denotes the total expense incurred in the test.
Model HZ-ORI HZ-DEST FSQ-TYK FSQ-Global Cost(\$)
ACC F1 ACC F1 ACC F1 ACC F1
LLM-Mob 74.03 0.7473 69.26 0.6901 16.2 0.1171 16.26 0.1066 43.8
MoBLLM 89.24 0.8921 80.15 0.7998 28.7 0.2528 25.42 0.2085 0.7
4.6 Adaptation under Different Prediction Situations
To investigate the adaptation under different situations (events, incidents), we further estimate the model prediction performance in the Hong Kong AFC data (mainly due to the wide data availability for the test). As shown in Table 5, the adaptation datasets involve individual trips occurring in metro network change (MNC), policy intervention (PI), special events (SE), and metro incidents (MI). Besides, each of the four datasets can be divided into 2 subsets according to origin prediction (ORI) and destination prediction (DEST). All the results of validating adaptability are shown in Table 6, where the results of HK-ORI and HK-DEST in regular scenarios are also listed for comparison.
Overall, the MoBLLM model has robust performance in different prediction situations and significantly outperforms all the baseline models. The deep learning based models have huge performance reduction in PI-ORI and PI-DEST. Especially in PI-DEST, their performance reduces by over 10% compared with the performance in HK-DEST, while the LLM-based models only have a margin reduction of about 2%. The probable reason is that the PI datasets are much different from the HK-ORI and HK-DEST used for training the deep learning models. Deep learning models have low transferability, which leads to significant performance variation in situations where the input data is inconsistent with the training data to some extent.
We further analyze the robustness of the MoBLLM model in each prediction situation. In datasets MNC and PI that have fewer stations than the training data, the MoBLLM model performance has an average decrease by about 7.5% on ACC and F1 for origin prediction compared to the performance in HK-ORI. The MoBLLM model performs worse in PI than MNC for both origin and destination prediction because there exist policy interventions of price discounts that probably cause a part of passengers to change their travel behaviors. In the datasets SE and MI that have special events and incidents influencing the demands, the MoBLLM model has a margin performance reduction of about 3% for the origin prediction. It is because a small number of passengers may change their travel choices. For destination prediction on these two datasets, the MoBLLM even has better performance than the one on HK-DEST, which implies that the special events may not significantly influence the destination of existing trips. Therefore, the MoBLLM model has a good generalization ability to adapt to different prediction situations covering metro network change, policy disturbances, special events, and incidents.
Table 5: The descriptions of adaptation datasets
Dataset User Number Test Samples Time Span Description
MNC 898 6318 2016/08-2016/11
Compared with training data,
it has less stations.
PI 844 5791 2014/07-2014/08
Compared with training data,
it has less stations and fare
discount in morning peak.
SE 1657 5010 2018/01-2018/03
Compared with training data,
it has the same number of stations
but all the samples are selected
from 35 days with live concerts.
MI 2919 4975 2018/01-2018/02
Compared with training data,
it has the same number of stations
but all the samples are selected
from 7 incident days with at least
10-minute delay.
Table 6: Model performance comparison under different scenarios.
Dataset DeepMove MobTCast MHSA LLM-Mob MoBLLM
HK-ORI ACC 82.81 83.28 83.55 62.81 87.86
F1 0.827 0.8324 0.8349 0.6494 0.8784
MNC-ORI ACC 74.39 78.8 78.69 63.49 83.1
F1 0.741 0.785 0.785 0.6412 0.83
PI-ORI ACC 71.52 75.91 76.51 58.97 79.4
F1 0.712 0.7552 0.7614 0.5911 0.7928
SE-ORI ACC 73.99 78.68 79.34 62.06 84.91
F1 0.7333 0.7798 0.7875 0.625 0.8482
MI-ORI ACC 75.47 80.78 81.56 65.33 85.59
F1 0.7497 0.8031 0.8129 0.6602 0.8554
HK-DEST ACC 65.45 65.49 66 54.23 70.93
F1 0.645 0.6478 0.6554 0.5334 0.7044
MNC-DEST ACC 55.22 59.43 58.76 55.67 71.85
F1 0.5416 0.586 0.583 0.5519 0.7229
PI-DEST ACC 51.54 55.08 55.15 53.76 68.71
F1 0.503 0.5436 0.5472 0.5322 0.6835
SE-DEST ACC 53.17 56.68 58.86 52.93 72.69
F1 0.5175 0.5569 0.5824 0.525 0.7235
MI-DEST ACC 60.7 63.41 64.78 61.95 79.92
F1 0.5951 0.6294 0.6436 0.6167 0.7971
5 Conclusions
This paper proposes the MoBLLM model for individual mobility prediction problem. It is obtained by adopting the PEFT method to tune an open-source LLM on the instruction dataset, which can be generated based on the real-world mobility data. Such fine-tuning enables the MoBLLM model to adapt to diverse individual prediction tasks across time scales, spatial scales, and conditional scenarios.
In the cases studies, six real-world mobility datasets from multiple cities are used to validate the performance of MoBLLM model. We systematically evaluate LoRA and its several advanced variants in fine-tuning to obtain the optimal MoBLLM model. Compared with the state-of-art baseline models on accuracy, the MoBLLM model averagely outperforms the best DL model (MHSA) by 50.44% and the LLM-Mob model based on commercial LLMs by 37.30%. Besides, the MoBLLM model also shows significant superiority in transferability and finance compared to the method based on commercial LLMs. In the adaptability evaluation on 4 trip datasets representing different conditional scenarios, the MoBLLM shows the robust performance and outperforms the best benchmark (MHSA) by 14.67% on average. These results of case studies comprehensively demonstrate the capacity of the MoBLLM model as the foundational individual mobility prediction model.
The MoBLLM model contributes to precise and robust provision of personalized information for transportation management, e.g., crowding and disruption management. However, note that the incident logs or event information are important for prediction. Future work will design the prompt (incorporating situational context) and explore the fine-tuning methods to further improve the MoBLLM’s adaptability for abnormal situations. Also, future validation of the MoBLLM model on data from more cities is valuable.
6 CRediT authorship contribution statement
Zhenlin Qin: Conceptualization, Data curation, Methodology, Visualization, Investigation, Formal analysis, Validation, Writing – original draft. Leizhen Wang: Methodology, Validation, Writing – review and editing. Yangcheng Ling: Data curation, Validation Zhenliang Ma: Conceptualization, Data curation, Methodology, Formal analysis, Writing – review and editing, Supervision. Francisco Câmara Pereira: Conceptualization, Formal analysis, Writing – review and editing, Supervision
Acknowledgements
The work was supported by the TRENoP center (Swedish Strategic Research Area in Transportation) at KTH Royal Institute of Technology, Sweden.
7 Data availability
The code, datasets (upon confidential check), experiments for this paper will be made public available at: https://github.com/qzl408011458/MoBLLM.
Appendix A
The following four instruction examples are respectively generated based on the instruction templates of corresponding base tasks defined in Section 3.1.
Figure A1: The generated instruction example for the task 1.
Figure A2: The generated instruction example for the task 2.
Figure A3: The generated instruction example for the task 3.
Figure A4: The generated instruction example for the task 4.

### v3

@@
 It is widely modeled by data-driven methods such as machine learning, deep learning, as well as classical econometric methods to capture key features of mobility patterns.
-However, such methods are hindered in promoting further transferability and generalization due to limited capacity to learn mobility patterns from different data sources, predict in out-of-distribution settings (a.k.a “zero-shot”).
-To address the challenge, this paper proposes the foundational MoBLLM model based on lightweight open-source large language models (LLMs) fine-tuned on mobility datasets.
-We conduct extensive experiments on six real-world mobility datasets to evaluate its performance and transferability for diverse prediction tasks across varying times (years), space (cities), and situations (e.g., disruptions).
-The MoBLLM model achieves best F1 and accuracy performance in all datasets Compared with the state-of-the-art deep learning models.
-It also shows better transferability and cost performance than commercial LLMs.
-Further experiment reveals its high adaptability to different situations under network change, policy intervention, special events, and incidents.
-These results demonstrate the foundational capability of the MoBLLM model in predicting individual mobility, which contributes to precise and robust provision of personalized information for transportation management.
+However, such methods are hindered in promoting further transferability and robustness due to limited capacity to learn mobility patterns from different data sources, predict in out-of-distribution settings (a.k.a “zero-shot”).
+To address this challenge, this paper introduces MoBLLM, a foundational model for individual mobility prediction that aims to learn a shared and transferable representation of mobility behavior across heterogeneous data sources.
+Based on a lightweight open-source large language model (LLM), MoBLLM employs Parameter-Efficient Fine-Tuning (PEFT) techniques to create a cost-effective training pipeline, avoiding the need for large-scale GPU clusters while maintaining strong performance.
+We conduct extensive experiments on six real-world mobility datasets to evaluate its accuracy, robustness, and transferability across varying temporal scales (years), spatial contexts (cities), and situational conditions (e.g., disruptions and interventions).
+MoBLLM achieves the best F1 score and accuracy across all datasets compared with state-of-the-art deep learning models and shows better transferability and cost efficiency than commercial LLMs.
+Further experiments reveal its robustness under network changes, policy interventions, special events, and incidents.
+These results indicate that MoBLLM provides a generalizable modeling foundation for individual mobility behavior, enabling more reliable and adaptive personalized information services for transportation management.
 keywords:
@@
 It facilitates a broad spectrum of applications in transportation domain, such as personalized services of POI or route recommendations and incentive mobility management (Ma and Zhang, 2022).
-Data-driven methods are dominant in this field including conventional Markov-based machine learning models (Gambs et al., 2012; Mathew et al., 2012; Lu et al., 2013) and deep learning (DL) models such as recurrent neural network (RNN) or Transformer (Feng et al., 2018; Kong and Wu, 2018; Xue et al., 2021).
+Data-driven methods are dominant in this field including conventional Markov-based machine learning models (2; W.
+Mathew, R.
+Raposo, and B.
+Martins (2012); X.
+Lu, E.
+Wetter, N.
+Bharti, A.
+J.
+Tatem, and L.
+Bengtsson (2013)) and deep learning (DL) models such as recurrent neural network (RNN) or Transformer (1; D.
+Kong and F.
+Wu (2018); H.
+Xue, F.
+Salim, Y.
+Ren, and N.
+Oliver (2021)).
 However, they are still limited to specialized training on the mobility dataset from a single city or data source, leading to low transferability of adapting to different travel contexts in different cities.
@@
 As fine-tuning techniques advance, LLMs can be applied to developing foundational models for other various disciplines (Wang et al., 2025; Awais et al., 2025).
-Inspired by the concept of foundational models in AI, we define the foundational mobility prediction model as one trained on extensive mobility data, capable of adapting to diverse prediction tasks across varying time scales (years), spatial scales (cities), and situational conditions (such as disruptions and interventions).
-LLMs have also inspired a range of applications in urban mobility (Ma et al., 2025; Wang et al., 2024), particularly in mobility prediction tasks.
+LLMs have also inspired a range of applications in urban mobility (Ma et al., 2025; Wang et al., 2024b), particularly in mobility prediction tasks.
 Specifically, typical LLM based methods like In-Context Learning (ICL) (Brown et al., 2020) and Chain-of-Thought (CoT) (Wang et al., 2023a; Wei et al., 2022) enable modeling mobility prediction problems with natural language (Wang et al., 2023b; Liang et al., 2024; Feng et al., 2024a).
-However, hallucination issues (Huang et al., 2025) may arise when applying these methods, as their use of commercial LLMs lacks specialized training on mobility datasets.
+However, hallucination issues may arise when applying these methods, as their use of commercial LLMs lacks specialized training on mobility datasets.
+The hallucination is first defined in natural language generation as generated content that is either nonsensical or unfaithful to the provided source content (Huang et al., 2025), which can be similarly defined as unexpected LLMs’ outputs under mobility prediction context.
 Therefore, the existing prompt engineering based methods are limited in generalizing prediction under different contexts (lack foundational capability).
@@
 They begin by aligning visited locations with specific intents, then leverage LLMs to predict the next location based on this enriched semantic context.
-From model validation perspectives, most of studies focus on prediction / accuracy but have not sufficiently demonstrated the LLMs’ capabilities of generalization and transferability across different data sources and prediction scenarios.
-To fill the gap, we propose a foundational individual mobility prediction model MoBLLM based on open-source LLMs.
-It uses the pre-trained LLMs as the core module with generic prediction prompts and introduces instruction tuning (Zhang et al., 2023c) to enhance the LLMs’ capacities in adapting to different prompt writing styles of users and ensuring high transferability on different types of mobility data (e.g., check-in data).
-It adopts the Parameter Efficient Fine-Tuning (PEFT) method to realize the parameter and cost efficiency of training compared with the conventional full-parameter fine-tuning method.
+From model validation perspectives, most prior studies focus primarily on prediction / accuracy but have not sufficiently demonstrated the LLMs’ capabilities of transferability and robustness across different data sources and prediction scenarios.
+Inspired by the notion of foundation models in artificial intelligence, we use the term foundational model in a transportation-science sense rather than in an architecture-centric AI sense.
+In this paper, a foundational mobility model refers to a reusable, general-purpose behavioral representation that supports multiple mobility-related prediction and policy-relevant tasks across heterogeneous cities, time periods, and contextual conditions.
+Unlike traditional transportation models that are calibrated separately for each dataset, city, or policy scenario, a foundational mobility model aims to learn a shared representation of individual mobility behavior that can be transferred, adapted, and reused across contexts.
+The novelty lies not in proposing a new neural architecture, but in establishing a unified modeling substrate for mobility prediction that remains valid across spatial scales (cities), temporal scales (years), and situational conditions (e.g., disruptions, pricing, and interventions).
+Under this definition, LLMs serve as a practical implementation mechanism for building such a foundation, because they provide a flexible representation of heterogeneous mobility data, traveler attributes, and contextual factors within a single transferable model.
 The key contributions include:
   • 1.
-    Propose a foundational, open-source LLM-based individual mobility prediction model, which can adapt to different data from multiple sources and cities, various individual mobility prediction tasks and even diverse prediction scenarios.
+    We introduce a general-purpose mobility foundation model for individual mobility prediction that provides a unified, transferable behavioral representation across heterogeneous data sources, cities, temporal periods, and situational contexts, enabling consistent modeling of mobility behavior beyond case-specific calibration.
   • 2.
-    Introduce an efficient instruction generation approach to produce plentiful CoT based prompts for adapting to different writing styles of users.
+    We develop a scalable instruction-based adaptation framework that enables large language models to be efficiently specialized for diverse mobility prediction tasks and user-specific data formats, supporting flexible transfer across datasets without retraining from scratch.
   • 3.
-    Validate the proposed method through extensive experiments on different mobility datasets, including GPS location, AFC, and check-in data from different cities over the world.
-It compares against state-of-the-art DL and LLM models in aspects of prediction accuracy and transferability (space, time, and situation).
+    We conduct extensive empirical evaluation on multi-city mobility datasets—including GPS trajectories, Automated Fare Collection (AFC) records, and check-in data—to systematically assess prediction accuracy, cross-city transferability, temporal generalization, and robustness to contextual variation, thereby establishing both the strengths and empirical boundaries of the proposed foundation model paradigm.
 The remainder of the paper is organized as follows: Section 2 reviews the recent approaches related to modeling individual mobility prediction in the literature.
@@
 The problem of individual mobility prediction consists of predicting an individual’s possible next location or trip information given their historical mobility sequence data (Ma and Zhang, 2022).
-In early stage, the Markov based methods are widely applied to modeling location transition probabilities, i.e, the next location prediction is based on several previous visited locations (Gambs et al., 2012; Mathew et al., 2012; Lu et al., 2013).
+In early stage, the Markov based methods are widely applied to modeling location transition probabilities, i.e, the next location prediction is based on several previous visited locations (2; W.
+Mathew, R.
+Raposo, and B.
+Martins (2012); X.
+Lu, E.
+Wetter, N.
+Bharti, A.
+J.
+Tatem, and L.
+Bengtsson (2013)).
 Although these methods are intuitive and interpretable, they have low performance due to the difficulty in learning complex mobility patterns.
@@
 (2016) models local temporal and spatial contexts by extending an recurrent neural network (RNN) layer with time-specific and distance-specific transition matrices.
-DeepMove (Feng et al., 2018) adopted a sequence to sequence (seq2seq)-based framework to capture the multi-level periodicity for mobility prediction from lengthy and sparse trajectories.
-The DeepTrip model (Zhang et al., 2023a) extended the DeepMove to predict the next trip information of the AFC metro data.
+DeepMove (1) adopted a sequence to sequence (seq2seq)-based framework to capture the multi-level periodicity for mobility prediction from lengthy and sparse trajectories.
+The DeepTrip model (Zhang et al., 2023b) extended the DeepMove to predict the next trip information of the AFC metro data.
 Other RNN-based models adopted the Long Short-Term Memory (LSTM) architecture to model the prediction problem (Kong and Wu, 2018; Li et al., 2020; Chen et al., 2020).
@@
 Similarly, LLM-MPE (Liang et al., 2024) predicted human mobility under public events by leveraging the unprecedented ability of LLM to process textual data.
-AgentMove (Feng et al., 2024a) achieved a generalized prediction of mobility for any city worldwide by introducing POI data from cities to capture the knowledge of the world.
-LingoTrip Qin et al.
-(2025) prompted LLMs to capture trip purposes, choice preferences and choice variability from individual’s historical trip data to comprehensively infer the next trip location.
+LingoTrip (Qin et al., 2025) prompted LLMs to capture trip purposes, choice preferences and choice variability from individual’s historical trip data to comprehensively infer the next trip location.
+Besides, agentic LLM frameworks integrating ICL, CoT and Retrieval-Augmented Generation (RAG) are applied to extending mobility prediction with more dimensions such as activity or intent inference.
+Specifically, AgentMove (Feng et al., 2024a) achieved a generalized prediction of mobility for any city worldwide by introducing POI data from cities to capture the knowledge of the world.
+ZHMF (Li et al., 2025) introduced a framework that integrates short-term trajectory records with structured long-term memory retrieved by RAG technology.
+LLMob (Wang et al., 2024a) adopted an agentic LLM framework that accounts for individual activity patterns and motivations.
 While these methods exhibit promising application potential, the performance still has significant room for improvement since they use commercial LLMs trained on millions of general corpora without training on any mobility dataset.
@@
 Although these methods enable LLMs to capture the mobility intents or preferences of individual by fine-tuning, they have not demonstrated the capability of adapting to different data sources across multiple cities and diverse prediction scenarios such as disruptions and interventions.
+The reviewed mobility-related works based on LLMs are summarized in Table 1 according to the primary modeling methods, data and problem settings.
 Overall, the existing methods based on prompt engineering or fine-tuning show high potential in modeling the individual mobility prediction problem and yield a wide range of valuable applications.
 However, they are still insufficient to build a foundational model for individual mobility prediction due to the unsolved issues in adapting to diverse prediction tasks across varying spatiotemporal scales and conditional scenarios.
+Table 1: Comparison between the LLM-based related Methods and MoBLLM.
+Model Main Method Data Type Foundational Capability Task
+LLM-Mob ICL and CoT Check-in, GPS trajectory Limited in accuracy Individual mobility prediction
+LLM-MPE ICL and CoT Taxi ridership Limited in accuracy
+Mobility prediction under public
+events
+LingoTrip ICL and CoT Metro trip Limited in accuracy Individual mobility prediction
+LLMob Agentic LLM Check-in Limited in accuracy Trajectory generation
+ZHMF Agentic LLM Check-in Limited in accuracy
+Intent-aware, cold-start mobility
+prediction
+AgentMove Agentic LLM Check-in, GPS trajectory Limited in accuracy Individual mobility prediction
+LIMP
+Agentic LLM,
+fine-tuning
+Call detail records Limited in data scope Intent-aware mobility prediction
+Mobility-LLM
+ICL and CoT,
+fine-tuning
+Check-in Limited in data scope Intent-aware mobility prediction
+MoBLLM
+ICL and CoT,
+fine-tuning
+Check-in, GPS trajectory,
+metro trip
+Strong Individual mobility prediction
 3 Methodology
@@
 We define end time $t_{e_{i}}$, and then the trip activity duration is $dur_{i}=t_{e_{i}}-l_{s_{i}}$.
-Task 4 (next trip destination prediction): Given a user’s sequence of $n-1$ chronological trips: $\textbf{Tr}=(Tr_{1},Tr_{2},\dots,Tr_{n-1})$, the task is to predict the next destination station of $Tr_{n}$.
+Task 4 (next trip destination prediction): Given a user’s sequence of n1n-1 chronological trips: $\textbf{Tr}=(Tr_{1},Tr_{2},\dots,Tr_{n-1})$, the task is to predict the next destination station of $Tr_{n}$.
 A trip $Tr_{i}$ includes start time of trip $t_{o_{i}}$, day of the week $w_{i}$, origin station $s_{o_{i}}$, duration $dur_{i}$ of last stay occurring before $Tr_{i}$, and destination station $d_{i}$.
@@
 The activity after the last trip of the day is not considered because it is exactly staying home the next day.
-The information of a trip mainly includes origin station $o$, destination station $d$, start time of trip $t_{o}$, end time of trip $t_{d}$, week of day $w$, which are commonly employed in trip prediction studies (Zhang et al., 2023b; Zhao et al., 2018).
+The information of a trip mainly includes origin station $o$, destination station $d$, start time of trip $t_{o}$, end time of trip $t_{d}$, week of day $w$, which are commonly employed in trip prediction studies (Zhang et al., 2023a; Zhao et al., 2018).
 However, we extend a trip with the duration of the last stay defined in Task 4, supposing it is associated with certain activity information.
@@
 Figure 2: The overview of the proposed MoBLLM framework.
+Given 4 base task prompt templates, Step 1 generates multi-style semi-complete instructions without any mobility data by prompting advanced commercial LLMs such as GPT-4o mini.
+Step 2 constructs the instruction dataset by randomly sampling and assembling real-world mobility data with the semi-complete instructions.
+Step 3 obtains the MoBLLM model (for prediction) by a standard supervised fine-tuning pipeline that uses PEFT approach to train an open-source LLM on the instruction dataset.
+Building upon the zero-shot prompting strategies of LLM-Mob (Wang et al., 2023b), which leveraged the general reasoning capabilities of commercial LLMs for mobility prediction, our work addresses a key limitation: these general-purpose models are not specialized for individual mobility, which constrains their predictive performance.
+While supervised fine-tuning methods such as PEFT provide a pathway to domain adaptation (Lu et al., 2025), they also risk catastrophic forgetting—where the model loses its ability to generalize to contexts beyond its fine-tuning dataset (Sun and Dredze, 2024).
+This raises concerns that fine-tuned LLMs may fail to transfer effectively to mobility data with distinct contexts (e.g., other cities or user populations) or structural types (e.g., different data formats or temporal resolutions).
+To mitigate this risk and ensure robust transferability, we construct the instruction dataset with two deliberate strategies.
+First, we incorporate multi-type mobility data from multiple cities, enabling the LLM to capture structural patterns that are invariant across heterogeneous sources.
+Second, we normalize all location labels in each dataset to a universal integer index (0 to M1M-1, where $M$ is the number of unique locations).
+This abstraction intentionally removes dataset-specific semantics, compelling the model to focus on the underlying mechanisms of mobility dynamics—such as temporal regularities, sequential dependencies, and routine stability—rather than memorizing spatial identifiers.
+In doing so, MoBLLM leverages the compositional and sequential reasoning capacities of LLMs to internalize a generalized representation of human movement behavior.
+Although this abstraction may reduce sensitivity to spatial semantics, it substantially enhances the model’s ability to generalize across datasets with different spatial and contextual properties, which is the defining property of a foundational mobility model.
+From a theoretical perspective, the generalization capability of MoBLLM can be interpreted through the lens of domain-invariant representation learning (Muandet et al., 2013; Arjovsky et al., 2019).
+By abstracting away spatial semantics through normalized indices, MoBLLM encourages the LLM to encode mobility sequences in a latent behavioral space that is less dependent on city-specific coordinate systems or naming conventions.
+In this space, prediction is guided primarily by learned transition dynamics and temporal regularities rather than explicit spatial memorization.
+This interpretation is conceptually aligned with ideas from causal and invariant representation learning (Schölkopf et al., 2021), in the sense that the model aims to capture behaviorally stable patterns—such as daily routines and temporal dependencies—that tend to persist across contexts.
+However, we emphasize that this remains a conceptual lens: no formal causal or invariance guarantees are claimed.
+Consequently, adaptation to unseen datasets is viewed as distributional alignment within this latent behavioral space, rather than proof of causal mechanism discovery.
 3.3 Prompt Template for Individual Mobility Prediction
@@
 Since most current LLMs are initially designed for dealing with various natural language tasks, there is a non-trivial mismatch between the original training objective of LLMs and our studied individual mobility prediction problem.
-To address the mismatch, we adopt the instruction tuning approach, which is commonly applied to enhance the capabilities and controllability of LLMs in other domain-specific tasks (Zhang et al., 2023c).
-The key to the approach is to construct an instruction-tuning dataset for LLMs’ fine-tuning.
+To address the mismatch, we adopt the instruction based fine-tuning approach, which is commonly applied to enhance the capabilities and controllability of LLMs in other domain-specific tasks (Zhang et al., 2023c).
+The key to the approach is to construct an instruction dataset for LLMs’ fine-tuning.
 Each sample in the dataset can be concisely represented as an $(Instruction,Output)$ pair.
@@
 Typically, distillation imparts knowledge and cognitive abilities from a competent teacher model to a simpler and more computationally efficient student model (Zhang et al., 2023c).
-Figure 4 shows the CoT-based prompt for generating multi-style instructions with three main parts (see Appendix for generated examples).
+Figure 4 shows the CoT-based prompt for generating multi-style instructions with three main parts (see A for generated examples).
 First, the LLMs should know how to generate diverse instructions related to time series prediction.
@@
 Then it combines with a piece of real-world mobility data from the database to get an $Instruction$.
-To align with the required prediction results of the task prompt, the corresponding $Output$ given an $Instruction$ is set as $\{prediction:place\_id\}$ where the $place\_id$ is the ground truth of the next location or origin / destination station defined in the Problem Definition.
+To align with the required prediction results of the task prompt, the corresponding $Output$ given an $Instruction$ is set as $\{prediction:place\_id\}$ where the place_idplace\_id is the ground truth of the next location or origin / destination station defined in the Problem Definition.
 Figure 3: The example of instruction data.
-For the data input, $<$history$>$ and $<$context$>$ are the user’s mobility sequences respectively, containing long-term and recent stays or trips.
-$<$target$>$ includes the information about current states except for location, e.g., the tap-out station and time of the last trip.
+For the data input, <<history>> and <<context>> are the user’s mobility sequences respectively, containing long-term and recent stays or trips.
+<<target>> includes the information about current states except for location, e.g., the tap-out station and time of the last trip.
 Figure 4: The prompt example to generate multi-style instructions. A prompt template of the base task is put at the end, prompting LLMs to learn the style.
@@
 In contrast, parameter-efficient fine-tuning (PEFT) methods significantly reduce memory requirements while maintaining high adaptation performance (Han et al., 2024).
-One of the most prominent PEFT techniques is Low-Rank Adaptation (LoRA) (Hu et al., 2021), which has been extensively used to fine-tune LLMs in mobility-related applications (Li et al., 2024; Gong et al., 2024).
+One of the most prominent PEFT techniques is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which has been extensively used to fine-tune LLMs in mobility-related applications (Li et al., 2024; Gong et al., 2024).
 Figure 5 shows that LoRA introduces two trainable matrices, $W_{1}\in\mathbb{R}^{d\times r}$ and $W_{2}\in\mathbb{R}^{r\times k}$, where the rank $r$ is significantly smaller than $\min(d,k)$.
@@
 $h=W_{0}X+\frac{\alpha}{r}W_{1}W_{2}X,$ (1)
-where $h$ denotes the output embedding aligned with mobility-related semantics in $X$. Here, $\alpha$ is a tunable scaling factor used to control the contribution of the low-rank update.
+where $h$ denotes the output embedding aligned with mobility-related semantics in $X$.
+Here, $\alpha$ is a tunable scaling factor used to control the contribution of the low-rank update.
 Instead of updating the full weight matrix, LoRA introduces a low-rank correction term $W_{1}W_{2}$ on top of $W_{0}$, allowing the model to efficiently adapt without retraining all parameters.
 Since the number of parameters in $W_{1}$ and $W_{2}$ is orders of magnitude smaller than in $W_{0}$, this approach significantly reduces GPU memory requirements.
-Empirical studies report that LoRA can achieve up to a 3 times reduction in memory usage while maintaining competitive performance (Hu et al., 2021), making it well-suited for fine-tuning LLMs on mobility datasets under limited resources.
+Empirical studies report that LoRA can achieve up to a 3 times reduction in memory usage while maintaining competitive performance (Hu et al., 2022), making it well-suited for fine-tuning LLMs on mobility datasets under limited resources.
 In the fine-tuning phase, the individual mobility prediction task is framed as a next-token generation task.
@@
 According to Tasks 3 and 4 for predicting origin and destination, each AFC data can be processed to obtain two subsets, such as HK-ORI and HK-DEST.
+The MoBLLM model is obtained by following the pipeline in Figure 2 that fine-tunes a pre-trained LLaMA-3.1-8B-Instruct  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct on the instruction dataset.
+We set MoBLLM’s temperature parameter to zero during inference to guarantee deterministic output across all case studies.
+In our transferability experiments, the training data were sampled from four datasets (Geolife, FSQ-NYC, HK-ORI and HK-DEST), while testing was conducted on four different datasets (FSQ-TYK, FSQ-Global, HZ-ORI, and HZ-DEST).
+As shown in Table 2, these datasets differ significantly in terms of data type (GPS trajectory, check-in, AFC), geographical context (Beijing, New York, Tokyo, Hong Kong, Hangzhou, and 415 global cities), and spatial coverage (ranging from dozens to over 100,000 unique locations).
+These differences suggest that the transfer experiments indeed involve substantial cross-dataset heterogeneity rather than trivial similarities.
 Table 2: The descriptions of datasets.
-Dataset Type User Number Test Samples District Task Type
-Geolife GPS trajectory 45 3459 Beijing task 1
-FSQ-NYC check-in 535 12141 New York task 2
-FSQ-TYK check-in 450 11250 Tokyo task 2
+Dataset Type Users Test Samples District Locations Time Span Task Type
+Geolife GPS trajectory 45 3459 Beijing 1145 2007/04-2012/08 task 1
+FSQ-NYC Check-in 535 12141 New York 3752 2011/10-2012/02 task 2
+FSQ-TYK Check-in 450 11250 Tokyo 15785 2012/04-2013/02 task 2
 FSQ-Global Check-in 1500 43500
@@
 countries.
-task 2
-HK-ORI AFC 500 5260 Hong Kong task 3
-HK-DEST AFC 500 5260 Hong Kong task 4
-HZ-ORI AFC 1200 15296 Hangzhou task 3
-HZ-DEST AFC 1200 15296 Hangzhou task 4
+112032 2012/04-2013/09 task 2
+HK-ORI AFC 500 5260 Hong Kong 92 2018/01-2018/03 task 3
+HK-DEST AFC 500 5260 Hong Kong 92 2018/01-2018/03 task 4
+HZ-ORI AFC 1101 15296 Hangzhou 81 2019/01-2019/02 task 3
+HZ-DEST AFC 1101 15296 Hangzhou 81 2019/01-2019/02 task 4
 In the case studies, we employ the following two metrics to evaluate the predictive performance of models:
@@
   • 2.
-    Weighted F1 measures the model’s performance in predicting frequently visited locations to avoid the issue of imbalance of individual’s visited locations, which is denoted as the abbreviation $F1$ in the following parts.
-    $F1=\frac{2TP}{2TP+FP+FN},$ (4)
-    where $TP$, $FP$ and $FN$ are respectively the number of true positive, false positive and false negative samples.
+    Weighted F1 measures the model’s performance in predicting frequently visited locations to avoid the issue of imbalance of individual’s visited locations.
+which is denoted as the abbreviation $F1$ in the following parts.
+    $F1=\sum_{i\in D}\frac{s_{i}}{S}\cdot\frac{2TP_{i}}{2TP_{i}+FP_{i}+FN_{i}},$ (4)
+    where $TP_{i}$, $FP_{i}$ and $FN_{i}$ are respectively the number of true positive, false positive and false negative samples given a ground truth location $l_{i}$.
+The $s_{i}$ represents samples of ground truth $l_{i}$ and the $S$ denotes the size of the dataset $D$.
+We use the F1 score weighted by the number of visits to emphasize the overall performance of the model in the most important locations.
 4.2 PEFT methods and parameters
@@
 Their specific values used in the PEFT methods in Table 3 are referred to in the corresponding literature, considering the trade-off between GPU memory, fine-tuning time, and model performance.
-All PEFT methods used in this research are based on a pre-trained LLaMA-3.1-8B-Instruct  [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct except the LoftQ starting with a 4-bit and 64-rank checkpoint of LLaMA-3-8B-Instruct  [^5] https://huggingface.co/LoftQ/Meta-Llama-3-8B-Instruct-4bit-64rank provided.
-About 100K training samples are used for fine-tuning and the batch size is set to 3 for all the methods.
+All PEFT methods used in this research are based on a pre-trained LLaMA-3.1-8B-Instruct except the LoftQ starting with a 4-bit and 64-rank checkpoint of LLaMA-3-8B-Instruct  [^5] https://huggingface.co/LoftQ/Meta-Llama-3-8B-Instruct-4bit-64rank provided.
+About 100K training samples are used for fine-tuning and the batch size is set to 3 for all the PEFT methods following common practices in LoRA-based fine-tuning studies that adopt very small batch sizes due to GPU memory constraints (Li et al., 2024).
 All the model training and experimental evaluations are conducted on a workstation with Intel(R) Xeon(R) w7-3465X CPU, 4 NVIDIA RTX 6000 Ada GPUs, and 512-Gb memory.
@@
 PiSSA 53.89 0.4829 35.21 0.3053 62.39 0.6259 41.77 0.3972 60 32
-LoftQ 58.5 0.537 40.92 0.3611 87.46 0.8745 70.48 0.6995 92.3 37.5
+LoftQ 58.5 0.537 40.92 0.3611 87.46 0.8745 70.48 0.6995 42.1 15.2
 LoRA+ 58.14 0.5346 39.86 0.3487 63.75 0.6348 41.52 0.3861 54.7 32
@@
   • 1.
-    DeepMove (Feng et al., 2018) uses recurrent networks with an attention mechanism to capture multi-scale temporal dependencies of mobility patterns.
+    DeepMove (1) uses recurrent networks with an attention mechanism to capture multi-scale temporal dependencies of mobility patterns.
   • 2.
@@
 The results are presented in Table 5, where the open-source MoBLLM model significantly outperforms the general-purpose commercial LLM.
+Both models perform worse on the check-in datasets than the AFC datasets since the former exhibit substantially higher sparsity in the set of possibly visited locations as shown in Table 2.
+This issue is particularly pronounced in FSQ-Global, which contains over 112,000 unique locations across 415 cities, resulting in extreme heterogeneity and long-tailed distributions.
+Such characteristics make prediction more difficult and contribute to the relatively low accuracy observed.
+Nevertheless, MoBLLM still demonstrates much stronger capabilities than LLM-Mob in capturing sparse mobility patterns, indicating that the main limitation arises from data sparsity and distribution complexity rather than the MoBLLM architecture itself.
 It demonstrates that an open-source LLM, i.e., LLaMA-3.1-8B-Instruct used in the case studies, can become an expert in different mobility prediction tasks by the proposed fine-tuning framework.
@@
 MoBLLM 89.24 0.8921 80.15 0.7998 28.7 0.2528 25.42 0.2085 0.7
-4.6 Adaptation under Different Prediction Situations
-To investigate the adaptation under different situations (events, incidents), we further estimate the model prediction performance in the Hong Kong AFC data (mainly due to the wide data availability for the test).
-As shown in Table 5, the adaptation datasets involve individual trips occurring in metro network change (MNC), policy intervention (PI), special events (SE), and metro incidents (MI).
-Besides, each of the four datasets can be divided into 2 subsets according to origin prediction (ORI) and destination prediction (DEST).
-All the results of validating adaptability are shown in Table 6, where the results of HK-ORI and HK-DEST in regular scenarios are also listed for comparison.
+4.6 Robustness under Different Prediction Situations
+[…]
 Overall, the MoBLLM model has robust performance in different prediction situations and significantly outperforms all the baseline models.
 The deep learning based models have huge performance reduction in PI-ORI and PI-DEST.
-Especially in PI-DEST, their performance reduces by over 10% compared with the performance in HK-DEST, while the LLM-based models only have a margin reduction of about 2%.
+Especially in PI-DEST, their performance reduces by over 10% compared with the performance in HK-DEST, while the LLM-based models only have a margin reduction of about 3%.
 The probable reason is that the PI datasets are much different from the HK-ORI and HK-DEST used for training the deep learning models.
@@
 We further analyze the robustness of the MoBLLM model in each prediction situation.
-In datasets MNC and PI that have fewer stations than the training data, the MoBLLM model performance has an average decrease by about 7.5% on ACC and F1 for origin prediction compared to the performance in HK-ORI.
+In datasets MNC and PI that have fewer stations than the training data, the MoBLLM model performance has an average decrease by about 4% on ACC for origin prediction compared to the performance in HK-ORI.
 The MoBLLM model performs worse in PI than MNC for both origin and destination prediction because there exist policy interventions of price discounts that probably cause a part of passengers to change their travel behaviors.
@@
 For destination prediction on these two datasets, the MoBLLM even has better performance than the one on HK-DEST, which implies that the special events may not significantly influence the destination of existing trips.
-Therefore, the MoBLLM model has a good generalization ability to adapt to different prediction situations covering metro network change, policy disturbances, special events, and incidents.
-Table 5: The descriptions of adaptation datasets
+Therefore, the MoBLLM model has a good robustness to adapt to different prediction situations covering metro network change, policy disturbances, special events, and incidents.
+Table 6: The descriptions of adaptation datasets.
 Dataset User Number Test Samples Time Span Description
@@
 10-minute delay.
-Table 6: Model performance comparison under different scenarios.
+[…]
 Dataset DeepMove MobTCast MHSA LLM-Mob MoBLLM
 HK-ORI ACC 82.81 83.28 83.55 62.81 87.86
-F1 0.827 0.8324 0.8349 0.6494 0.8784
-MNC-ORI ACC 74.39 78.8 78.69 63.49 83.1
-F1 0.741 0.785 0.785 0.6412 0.83
-PI-ORI ACC 71.52 75.91 76.51 58.97 79.4
-F1 0.712 0.7552 0.7614 0.5911 0.7928
+[…]
+MNC-ORI ACC 74.39 78.80 78.69 63.49 83.10
+[…]
+PI-ORI ACC 71.52 75.91 76.51 58.97 79.40
+[…]
 SE-ORI ACC 73.99 78.68 79.34 62.06 84.91
-F1 0.7333 0.7798 0.7875 0.625 0.8482
+[…]
 MI-ORI ACC 75.47 80.78 81.56 65.33 85.59
-F1 0.7497 0.8031 0.8129 0.6602 0.8554
-HK-DEST ACC 65.45 65.49 66 54.23 70.93
-F1 0.645 0.6478 0.6554 0.5334 0.7044
-MNC-DEST ACC 55.22 59.43 58.76 55.67 71.85
-F1 0.5416 0.586 0.583 0.5519 0.7229
+[…]
+HK-DEST ACC 65.45 65.49 66.00 54.23 70.93
+[…]
+MNC-DEST ACC 55.22 59.43 58.76 55.67 72.29
+[…]
 PI-DEST ACC 51.54 55.08 55.15 53.76 68.71
-F1 0.503 0.5436 0.5472 0.5322 0.6835
+[…]
 SE-DEST ACC 53.17 56.68 58.86 52.93 72.69
-F1 0.5175 0.5569 0.5824 0.525 0.7235
-MI-DEST ACC 60.7 63.41 64.78 61.95 79.92
-F1 0.5951 0.6294 0.6436 0.6167 0.7971
+[…]
+MI-DEST ACC 60.70 63.41 64.78 61.95 79.92
+[…]
 5 Conclusions
@@
 Besides, the MoBLLM model also shows significant superiority in transferability and finance compared to the method based on commercial LLMs.
-In the adaptability evaluation on 4 trip datasets representing different conditional scenarios, the MoBLLM shows the robust performance and outperforms the best benchmark (MHSA) by 14.67% on average.
+In the robustness evaluation on 4 trip datasets representing different conditional scenarios, the MoBLLM shows the robust performance and outperforms the best benchmark (MHSA) by 14.67% on average.
 These results of case studies comprehensively demonstrate the capacity of the MoBLLM model as the foundational individual mobility prediction model.
+This work has explored the application of Zero-Shot CoT prompting on the fine-tuned LLMs for individual mobility prediction.
+It is important to note, however, that the sequential nature of CoT reasoning introduces a vulnerability to error propagation.
+An incorrect fact extraction or flawed logical inference in an early step can cascade, adversely affecting all subsequent steps and the final outcome.
+This lack of fault tolerance presents a challenge for the deployment of fully autonomous reasoning systems.
+Future research should focus on integrating robustness mechanisms such as self-consistency or verification prompting.
 The MoBLLM model contributes to precise and robust provision of personalized information for transportation management, e.g., crowding and disruption management.
 However, note that the incident logs or event information are important for prediction.
-Future work will design the prompt (incorporating situational context) and explore the fine-tuning methods to further improve the MoBLLM’s adaptability for abnormal situations.
-Also, future validation of the MoBLLM model on data from more cities is valuable.
+Future work will involve the design of context-aware prompting strategies and the exploration of fine-tuning techniques to enhance MoBLLM’s adaptability, particularly in abnormal situations.
+Another promising direction involves enhancing the interpretability of LLM-generated predictions to gain deeper insights into the mechanisms governing human mobility behaviors.
+We also plan to extend and redevelop the proposed foundational model by evaluating its transferability to a wider spectrum of mobility prediction tasks, such as trip duration estimation.
+Additionally, we will explore practical applications of MoBLLM in personalized mobility research, including trip pattern analysis, to demonstrate its broader utility.
 6 CRediT authorship contribution statement
@@
 The code, datasets (upon confidential check), experiments for this paper will be made public available at: https://github.com/qzl408011458/MoBLLM.
-Appendix A
+Appendix A Instruction Examples
 The following four instruction examples are respectively generated based on the instruction templates of corresponding base tasks defined in Section 3.1.
@@
 Figure A4: The generated instruction example for the task 4.
+Appendix B Hallucination Cases
+For the base LLaMA-3.1-8B-Instruct model, we test it on the datasets used in the case studies using the same prompts as the MoBLLM model.
+As shown in Figures B1 andB2, the base model often produces irrelevant content, such as Python code snippets unrelated to the prediction task, which we identify as hallucinations in this context.
+Figure B1: The hallucination example for the task 1.
+Figure B2: The hallucination example for the task 3.
+Appendix C Hyperparameter Settings of PEFT
+For PEFT methods tested in Section 4.2, we present the settings of essential hyperparameter scaling factor $\alpha$ and rank $r$ as shown in Table 1.
+Table 1: The hyperparameter settings of PEFT methods.
+PEFT Scaling Factor Rank
+LoRA 64 64
+OLoRA 16 64
+EVA 1 16
+PiSSA 64 64
+LoftQ 16 64
+LoRA+ 8 32
+rsLoRA 8 32
+QLoRA 8 32
+Appendix D Ablation Study on Instruction Training Data
+In this study, we fine-tune the LLaMA-3.1-8B-Instruct model using single-style instruction data, while keeping all other experimental conditions consistent with those of MoBLLM.
+We referred to this variant as MoBLLM-S.
+The training set incorporates the same mobility data as MoBLLM but is constructed exclusively with predefined base task prompt templates, without introducing stylistic diversity.
+To assess the model’s robustness to diverse user writing styles, input prompts are generated following an approach analogous to that outlined in Section 3.4.
+Unlike the single-style input prompts aligned with LLM-Mob used in Section 4, the prompts in this experiment are generated by DeepSeek-Chat with a temperature of 0.3 based on the base task prompt templates to simulate a variety of writing styles.
+Note that employing a different LLM (DeepSeek-Chat) for generating test cases prevents leakage of styles identical to those produced by GPT-4o mini (generating training instructions for MoBLLM), thereby ensuring a fair comparison.
+The results of MoBLLM-S and MoBLLM are presented in Table 1, which demonstrate the advantages of multi-style instructions for fine-tuning to achieve higher model performance.
+Table 1: Ablation study on instruction training data
+Model Geolife FSQ-NYC HK-ORI HK-DEST
+ACC F1 ACC F1 ACC F1 ACC F1
+MoBLLM-S 57.79 0.5318 40.98 0.3630 83.71 0.8362 70.46 0.6991
+MoBLLM 58.37 0.5381 41.14 0.3638 84.26 0.8421 71.44 0.7093
+Appendix E Model Performance for In-/Frequent Visited Locations
+This study further investigates model performance on visited locations with different frequencies.
+First, we define the visited frequency as the ratio of the target location (ground truth) appearing in the model input (mobility sequence).
+To better present the stratified results, we select Geolife, FSQ-NYC and FSQ-TYK datasets and split their samples into 4-level visited frequencies as shown in Table 1.
+The results of Table 2 show that both LLM-Mob and MoBLLM models perform well in high-frequency samples but hardly make predictions on the target locations that have never been seen in the model input.
+Compared with the commercial LLM for general purpose, MoBLLM performs significantly better in the low- to medium- frequency range from 0 to 0.5, implying that fine-tuning can effectively improve the LLMs’ performance in predicting samples with such visited frequencies.
+Table 1: The sample proportions (%) of 4-level visited frequencies.
+Frequency Geolife FSQ-NYC FSQ-TYK
+0 12.44 1.61 1.07
+(0, 0.2] 32.58 66.77 86.89
+(0.2, 0.5] 44.72 27.22 11.52
+(0.5, 1] 10.26 4.4 0.52
+Table 2: The performance comparison (ACC) under different visited frequencies.
+Frequency Geolife FSQ-NYC FSQ-TYK
+LLM-Mob MoBLLM LLM-Mob MoBLLM LLM-Mob MoBLLM
+0 0 0.93 0 0 0 0
+(0, 0.2] 29.9 44.63 12.06 25.27 10.24 22.09
+(0.2, 0.5] 66.77 77.38 61.27 73.68 59.03 78.16
+(0.5, 1] 90.14 87.61 90.24 92.68 94.92 96.61